OpenStudy (anonymous):

I'll Give A medal to best answer Evaluate g(x) for any value within the domain. Interpret the meaning of your solution in the context of your problem

OpenStudy (anonymous):

{(0.5,64) , (1,128) , (2, 256)}

OpenStudy (agent0smith):

Use any two of the points to make the equation of a line... first find the slope, \[\large m = \frac{ y_2-y_1 }{ x_2 - x_1 }\]

OpenStudy (anonymous):

256-128= 128 2-1= 1 m=128/1

OpenStudy (agent0smith):

Now use that m and any one of the points in: \[\large y-y_1 = m(x-x_1)\]

OpenStudy (anonymous):

y-128=128/1(x-256) ?

OpenStudy (agent0smith):

yes but there's no need for 128/1... 128/1 is 128

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

so that's my answer?

OpenStudy (agent0smith):

No, simplify it first

OpenStudy (agent0smith):

y-128=128/1(x-256) ? and 256 is not an x-value, it's a y-value...

OpenStudy (anonymous):

y=128x+32896 ?

OpenStudy (anonymous):

It sounds a bit bizarre, i think i did it wrong

OpenStudy (agent0smith):

see my last post\[\large y-y_1 = m(x-x_1) \] (x1, y1) not two y values as you used

OpenStudy (anonymous):

y=128x

OpenStudy (anonymous):

is that my answer?

OpenStudy (agent0smith):

Yes

OpenStudy (anonymous):

thanks